(b) By considering the derivative of the function f : [−1, 1] → R defined by f(x) = \frac{2x}{\:x^2+1} , show that f^{-1} exists and find \left(f^{-1}\right)^' ( \frac{4}{\:5} ).


1
Expert's answer
2021-05-27T16:13:43-0400

The inverse function theorem helps us to validate that "f^{-1}" exist.

And, we have that

"(f^{-1})'(x) = \dfrac{1}{f'(f^{-1}(x))}"

First, we find the derivative of f using quotient rule

"f'(x) = \dfrac{(x^2+1)(2)-(2x)(2x)}{(x^2+1)^2} = \dfrac{2(1-x^2)}{(x^2+1)^2}"

Next, we find the inverse of f.

Suppose y=f(x) for x in the defined domain and y in R. So, we have that

"y= \dfrac{2x}{x^2+1} \\ x^2y+y=2x\\x^2y-2x=-y\\ x^2-\frac{2x}{y}=-1\\(x-\frac{1}{y})^2=\dfrac{1-y^2}{y^2}\\ x-\frac{1}{y} = \pm \sqrt{\dfrac{1-y^2}{y^2}} \\ x= \dfrac{1 \pm \sqrt{1-y^2}}{y}"

By putting x everywhere, we see y, we know that the inverse function is

"f^{-1}(x)= \dfrac{1 \pm \sqrt{1-x^2}}{x}"

Taking positive square roots, we have that

"f^{-1}(x)= \dfrac{1 + \sqrt{1-x^2}}{x}"

So,

"(f^{-1})'(\frac{4}{5}) = \dfrac{1}{f'(f^{-1}(\frac{4}{5}))}"

Consider,

"f^{-1}(\frac{4}{5}) = \dfrac{1 + \sqrt{1-(\frac{4}{5})^2}}{\frac{4}{5}} = 2"

So, "(f^{-1})'(\frac{4}{5}) = \dfrac{1}{f'(f^{-1}(\frac{4}{5}))}" becomes "(f^{-1})'(\frac{4}{5}) = \dfrac{1}{f'(2)}"

Also, "f'(2) = \dfrac{2(1-2^2)}{(2^2+1)^2} = -\dfrac{6}{25}"

Hence, "(f^{-1})'(\frac{4}{5}) = -\dfrac{25}{6}"